OpenStudy (anonymous):

g(x)=(1-2x)^2(x-3), g'(x)= (1-2x)^2-4(1-2x)(x-3), what is g''(x) equalls to?

OpenStudy (anonymous):

g''x = 24x - 32

OpenStudy (anonymous):

\[g'(x)=1-4x-4x^2+4x+12+8x^2-24x=12x^2-32x+1\]\[g''(x)=24x-32\]

OpenStudy (anonymous):

what are the y and x intercepts of g(x)=(1-2x)^2(x-3)?

OpenStudy (anonymous):

To find the y-intercept, set replace x with 0 in the equation, then solve for g(x) To find the x-intercept, replace g(x) with 0 in the equation then solve for x.

OpenStudy (anonymous):

how can i find where the function is increasing and decreasing?

OpenStudy (anonymous):

You have to graph it

OpenStudy (anonymous):

i need to show my calculations